# The AI_INFN Platform: Artificial Intelligence Development in the Cloud

### v1

@@ Abstract
Machine Learning (ML) is driving a revolution in the way scientists design, develop, and deploy data-intensive software. However, the adoption of ML presents new challenges for the computing infrastructure, particularly in terms of provisioning and orchestrating access to hardware accelerators for development, testing, and production. The INFN-funded project $\mathrm{AI\_INFN}$ (Artificial Intelligence at INFN) aims at fostering the adoption of ML techniques within INFN use cases by providing support on multiple aspects, including the provisioning of AI-tailored computing resources. It leverages cloud-native solutions in the context of INFN Cloud, to share hardware accelerators as effectively as possible, ensuring the diversity of the Institute’s research activities is not compromised. In this contribution, we provide an update on the commissioning of a Kubernetes platform designed to ease the development of GPU-powered data analysis workflows and their scalability on heterogeneous distributed computing resources, also using the offloading mechanism with Virtual Kubelet and InterLink API. This setup can manage workflows across different resource providers, including sites of the Worldwide LHC Computing Grid and super computers as CINECA Leonardo, providing a model for use cases requiring dedicated infrastructures for different parts of the workload. Initial test results, emerging case studies and integration scenarios will be presented with functional tests and benchmarks.
@@ 1 Introduction
The rapid proliferation of Machine Learning (ML) and Artificial Intelligence (AI) is fundamentally reshaping the landscape of data-intensive scientific research. In fields such as High Energy Physics (HEP), where experiments generate vast and complex datasets, these advanced computational techniques are becoming indispensable for tasks ranging from data acquisition and simulation to final physics analysis [1]. The Italian National Institute for Nuclear Physics (INFN), a leader in subnuclear, nuclear, and astroparticle physics, recognizes this paradigm shift. In response to these evolving computational demands, INFN is restructuring its computing infrastructure to adapt to advances in AI and Cloud technologies, enhancing resources [2] and expanding services through INFN Cloud [3, 4]. Within this context, INFN has launched the $\mathrm{AI\_INFN}$ initiative, a strategic effort to integrate AI-powered technologies across its research domains. The core mission of the $\mathrm{AI\_INFN}$ initiative is to create a cohesive ecosystem that connects infrastructure, algorithms, and applications, thereby fostering collaboration and accelerating discovery. This mission is structured around four primary work packages:
@@
    Infrastructure Provisioning: Procuring and operating a state-of-the-art computing infrastructure equipped with hardware accelerators like GPUs.
@@
    Training and Education: Organizing training events, such as advanced hackathons [5], to promote the adoption of modern ML techniques.
@@
    Community Building: Cultivating a community of ML experts and developers across the various INFN units to share knowledge and best practices.
@@
    Future Technologies: Developing expertise with next-generation hardware accelerators beyond GPUs, including Field-Programmable Gate Arrays (FPGAs) and Quantum Processors, for AI model training and inference.
The $\mathrm{AI\_INFN}$ platform [6, 7] is a cloud-native toolset developed in collaboration with the $\mathit{DataCloud}$ initiative, operating $\mathrm{INFN\ Cloud}$, to support the activities of these goals. It is a cloud-native, scalable, and user-centric environment designed to empower INFN researchers with the tools they need to pioneer the next generation of scientific breakthroughs.

### v2

@@ Abstract
Machine Learning (ML) is profoundly reshaping the way researchers create, implement, and operate data-intensive software. Its adoption, however, introduces notable challenges for computing infrastructures, particularly when it comes to coordinating access to hardware accelerators across development, testing, and production environments. The INFN initiative $\mathrm{AI\_INFN}$ (Artificial Intelligence at INFN) seeks to promote the use of ML methods across various INFN research scenarios by offering comprehensive technical support, including access to AI-focused computational resources. Leveraging the INFN Cloud ecosystem and cloud-native technologies, the project emphasizes efficient sharing of accelerator hardware while maintaining the breadth of the Institute’s research activities. This contribution describes the deployment and commissioning of a Kubernetes-based platform designed to simplify GPU-powered data analysis workflows and enable their scalable execution on heterogeneous distributed resources. By integrating offloading mechanisms through Virtual Kubelet and the InterLink API, the platform allows workflows to span multiple resource providers, from Worldwide LHC Computing Grid sites to high-performance computing centers like CINECA Leonardo. We will present preliminary benchmarks, functional tests, and case studies, demonstrating both performance and integration outcomes.
@@ 1 Introduction
The rapid expansion of Machine Learning (ML) and Artificial Intelligence (AI) is fundamentally transforming data-intensive scientific research. In domains like High Energy Physics (HEP), where experiments produce massive and intricate datasets, these advanced computational methods have become essential for tasks ranging from data collection and simulation to final physics analyses [1]. The Italian National Institute for Nuclear Physics (INFN), a leading institution in subnuclear, nuclear, and astroparticle physics, acknowledges this shift and is adapting its computing infrastructure to meet the growing demands of AI and cloud technologies. Through INFN Cloud [2, 3], resources [4] are being upgraded and services broadened to support this transition. Within this framework, the $\mathrm{AI\_INFN}$ initiative has been launched to integrate AI-driven technologies across INFN’s research areas. Its primary goal is to build a unified ecosystem connecting infrastructure, algorithms, and applications, fostering collaboration, and accelerating scientific discoveries. The initiative is structured around four main work packages:
@@
    Infrastructure Provisioning: Establishing and operating advanced computing infrastructure equipped with GPUs and other hardware accelerators.
@@
    Training and Education: Hosting workshops, hackathons, and training sessions to promote proficiency in modern ML techniques.
@@
    Community Building: Developing a network of ML practitioners and developers across INFN units to facilitate knowledge sharing and best practices.
@@
    Future Technologies: Exploring next-generation hardware for AI applications, including FPGAs and quantum processors, to enhance model training and inference.
The $\mathrm{AI\_INFN}$ platform [5, 6], developed in partnership with the $\mathit{DataCloud}$ initiative, operating $\mathrm{INFN\ Cloud}$, provides a cloud-native, scalable, and user-focused environment. It empowers INFN researchers with the necessary tools to advance AI-powered scientific exploration and support the development of innovative workflows and analyses.